A manufacturer of Christmas decorations has in 20x8 produced 1,000,000 large Christmas balls and 2,000,000 small
Christmas balls. Both products use the same raw materials and the production process
requires only minor adjustments when changing product. The company uses equivalence calculation for
to calculate the cost of the products.
The Christmas decoration manufacturer has calculated that a small Christmas ball requires 40% of the raw materials that go with a
large Christmas ball, 60% of direct salary and 50% of the indirect manufacturing costs involved in
production of large Christmas balls. The supplement for sales and administration is calculated as a percentage supplement
on manufacturing cost.
Total costs in 20x8 were:

Direct materials NOK 450,000
Direct salary NOK 660,000
Indirect manufacturing costs NOK 900,000
Sales and administration costs NOK 600,000

What is the cost of a small Christmas ball?

a. 0,51 NOK
b. 0,66 NOK
c. 0,87 NOK
d. 1,00 NOK
Answer #1

Answer : b. 0.66NOK

Calculation :

1. Direct material = 450,000*40% = NOK 180,000

2. Direct salary = 660,000*60% = NOK 396,000

3. Indirect manufacturing cost = 900,000*50%

= NOK 450,000

4. sales &administration cost= same percentage of indirect manufacturing cost = 600,000*50%

= NOK 300,000

Total cost = 180,000+396,000+450,000+300,000

= NOK 1,326,000

Number of small Christmas ball produced

= 2,000,000 balls

Cost per small Christmas ball = 1,326,000/2,000,000 = NOK 0.66